Home‎ > ‎Reports‎ > ‎Crew104‎ > ‎

day06


Crew 104 Log Book for April 28, 2011

Photos of the day





Commander's Check-In Report

Jon Clarke Reporting
 
Crew Physical Status: Good

Time Departed/Returned from EVA:  Excursion by six crew members to Green River and Horseshoe canyon while Sarah Thompson trouble shot the MARTE drill software in the Hab. Departed 9:30 am, returned 9:30 pm.

Brief Narrative of Field Mission Results: Observed Green River exhumed and inverted channels in the Ceadar Mt and Saltwash units, these were identified as Mars analogues by Becky Williams of the PSI as Mars analogues.  Endolith and chasmolith samples were collected from the silicified sandstones at the top of the Cedar Mt inverted channels. Horseshoe canyon was also visited and the crew familiarized themselves with the spectacular aeolain sandstones of the Navajo Fm, and the characteristics of aeolian sandstones in general.

EVA Data/Interpretations:  Pending analysis

Engineering/Hab Maintenance: No issues other than the MARTE drill.

Report Transmission Schedule: We will send the engineering report in the next hour

Plans for Tomorrow: More trouble shooting on MARTE, specially of mechanical issues.  Ongoing lab work on bulk organic and DNA extractions of surface and subsurface samples, catching up on reporting.

Inventory: No issues

Support Requested:  Any mousetraps in the Hab?  There have been two mouse sightings.

Miscellaneous: None

Engineering Report

David Willson Reporting

Generator/Electricity:

Kitty (Used? Yes/No): Used Continuously and not stopped

Honey (Diesel Generator) (Used? Yes/No): Not Used

Wendy (Gas Generator) (Used? Yes/No): Not Used

Battery Charge Level: Not Checked

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Generator operating very well continuously and not stopped


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 25% in Tank

Propane Tank (Full/0.75/0.5/0.25/Empty): 32.5%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): N/A

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Full

Honey Oil Quantity (# of Quarts in storage): Still Very Low, Will be toped up tomorrow

ATV Oil Quantity (# of Quarts in storage): Not Checked

Car Oil Quantity (# of Quarts in storage): Not Checked


Water Status:

Outside Potable Water Tank Level (inches from bottom): 85% Full

Trailer Potable Water Tank Level (inches from bottom): 15 cm

Hab Potable Water Tank Level (inches from bottom): 52%

Potable Water Meter Reading: 457134

In to GreenHab Meter Reading: N/A

Notes/Comments: Green Hab not operational but water is being recycled


GreenHab:

Condition of plants in Tank 1: 5%

Condition of plants in Tank 2: Full

Condition of plants in Tank 3: 15% Full

Greenhab Notes/Comments: Green house not used


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL): Operating

Used (yes/no): NO

Oil Checked (yes/no):

Tire Status:

Odometer at end of day:

Notes/Comments on Hab Car:


ATV 1 (Opportunity): Operating

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 2 (Spirit):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 3 (Viking I):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 4 (Viking II):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): A/C not operating 69 degrees F

Actual temperature upstairs (Farenheit): 69 degrees F

Thermostat setting downstairs (Farenheit): 65 degrees F

Actual temperature downstairs (Farenheit):


Computers/Networking Infrastructure: Operating

Notes/comments on internet/computers?: All OK


General Engineering Notes/Comments: